Suppose that 5 men in 100 are colorblind, while 24 women in 10,000 are colorblind. Compute the probability that a colorblind individual is male. Assume that the population contains an equal number of males and females. Round your answer to two decimal places. (If necessary, consult a list of formulas.)
Answer #1

Let,

M: Male individual

F: Female individual

C: Colorblind Individual

Given:

P(M) = P(F) = 1/2-0,5

PlClM) 5/100= 0.05

PlCF) 24/10000-0.0024

To find:

P(M|C)=?

By Bayes formula:

PAB) = P(BIA) * P(A)

thus,

P(CM) PM) P(C) P(MIC).............(1)

Also by total probability theorem,

P(A)= sum P(E_i) * P(A|E_i)

thus,

P(C) = P(M) * P(CIM) P(F) * P(CF)

P(C) (0.5 * 0.05) (0.5 * 0.0024)

P(C) 0.0262

Now in (1):

0.05 0.5 =-0.0262 P(MIC) _ 0.05 *

Pl MIC) 0.95419847328 ~ 0.95

Thus the probability that a colorblind individual is blind is 0.95.

Please upvote. Thank you.